OpenStudy (anonymous):

find (f o g)(x) find the domain of the composite function: f(x)=13/x+2; g(x)= x-5

OpenStudy (anonymous):

\[f(g(x))=f(x-5)=\frac{13}{(x-5)+2}=\frac{13}{x-3}\]

OpenStudy (anonymous):

i just replaced the x in f(x) by x - 5

OpenStudy (anonymous):

of course you cannot divide by 0, so the domain is all numbers except where x - 3 = 0 i.e all numbers except x = 3